Alliance, Inc. sells gas lamps to consumers through retail outlets. Total industry sales for Alliance's relevant market last year were $100 million, with Alliance's sales representing 5% of that total. Contribution margin is 25%. Alliance's sales force calls on retail outlets, and each sales rep earns $50,000 per year plus 1% commission on all sales. Retailers receive a 40% margin on selling price and generate average revenue of $10,000 per outlet for Alliance.

a. The marketing manager has suggested increasing consumer advertising by $200,000. By how much would dollar sales need to increase to break even on this expenditure? What increase in overall market shares does this represent?

b. another suggestion is to hire two more sales reps to gain new consumer retail accounts. How many new retail outlets would be necessary to break even on the increased cost of adding 2 sales reps?

c. A final suggestion is to make a 10% across the board price reduction. By how much would dollar sales need to increase to maintain Alliance's current contribution?

d. Which suggestion do you think Alliance should implement? Explain your recommendation.
Answer #1

a. Existing Market Share = 5%

Advertisement Expenditure = $200000

Overall Contribution Margin = 25%

Required Increase in sales to break even on Advertisement Expenditure = $200000/25%

   =$800000

Increase in overall Market share = 800000/100000000*100

=.80%

b. Retail Outlet Contribution Margin = 40%

Fixed Salary of Sales Representative of Retail Outlet = $50000 * 2

= $100000

Variable Salary of Sales Representative of Retail Outlet = 1% of Sales

Variable Salary i.e. Commission on sales is need not to be considered to calculate the break even as it has been already considered in the Contribution Margin.

Required sales for 2 new outlet to break even = $100000/40%

= $250000

Note: Since one Sales representative is required for each Retail outlet, hence we assume that Alliance Inc. has open 2 more Retail Outlets

c. Existing Sales of Alliance Inc. = $100,000,000 * 5%

= $5,000,000

Existing Contribution = $5,000,000 * 25%

= $ 1,250,000

Proposed Reduction in sales price = 10%

Proposed Contribution Margin = 25 - 10

=15%

Required sales to maintain the existing contribution level= $1,250,000/15%

= $8,333,333 (Rounded off)

Required increase in Sales = 8,333,3333 - 5,000,0000

= $3,333,333